Documenta Mathematica

Band 21, 2016

Sebastian Herpel and David I. Stewart On the Smoothness of Normalisers, the Subalgebra Structure of Modular Lie Algebras,

and the Cohomology of Small Representations  1–37

Baptiste Morin Milne’s Correcting Factor

and Derived De Rham Cohomology  39–48

Amalendu Krishna and Jinhyun Park

On Additive Higher Chow Groups of Affine Schemes  49–89

Thomas Geisser and Alexander Schmidt Tame Class Field Theory for Singular Varieties

over Algebraically Closed Fields  91–123

Shuichi Hayashida Lifting from two Elliptic Modular Forms to Siegel Modular Forms of Half-Integral Weight

of even Degree  125–196

Michael Lönne and Matteo Penegini On Asymptotic Bounds for the Number of Irreducible Components of the Moduli Space of

Surfaces of General Type II  197–204

Aurélien Djament, Teimuraz Pirashvili et Christine Vespa Cohomologie des foncteurs polynomiaux

sur les groupes libres  205–222

Bryden Cais, Tong Liu

On F -Crystalline Representations  223–270

A. Campillo, F. Delgado, and S. M. Gusein-Zade Equivariant Poincaré Series

and Topology of Valuations  271–286

Frank Gounelas

Free Curves on Varieties  287–308

Søren Eilers, Xin Li, and Efren Ruiz The Isomorphism Problem for Semigroup C^{∗}-Algebras

of Right-Angled Artin Monoids  309–343

Kam-Fai Tam Some Endoscopic Properties of The Essentially Tame Jacquet-Langlands

Correspondence  345–389

K. Yajima Remarks on L^{p}-Boundedness of Wave Operators for

Schrödinger Operators with Threshold Singularities  391–443

Giulio Codogni Hyperelliptic Schottky Problem

and Stable Modular Forms  445–466

Brooks Roberts and Ralf Schmidt

Some Results on Bessel Functionals for GSp(4)  467–553

David Burns, Masato Kurihara, and Takamichi Sano

On Zeta Elements for 𝔾_{m}  555–626

B. Krötz, E. Sayag, H. Schlichtkrull The Harmonic Analysis of Lattice Counting

on Real Spherical Spaces  627–660

Kieran G. O’Grady

Decomposable Cycles and Noether-Lefschetz Loci  661–687

Holger Deppe p-adic L-Functions of Automorphic Forms

and Exceptional Zeros  689–734

Pavle V. M. Blagojević, Florian Frick, Albert Haase, Günter M. Ziegler Hyperplane Mass Partitions

via Relative Equivariant Obstruction Theory  735–771

Zur Izhakian, Manfred Knebusch, Louis Rowen Quadratic and Symmetric Bilinear Forms

on Modules with Unique Base over a Semiring  773–808

Charles Starling C*-Algebras of Boolean Inverse Monoids –

Traces and Invariant Means  809–840

Michael Anshelevich, John D. Williams Operator-Valued Monotone Convolution Semigroups

and an Extension of the Bercovici-Pata Bijection.  841–871

Steffen Kionke, Jean Raimbault With an Appendix by Nathan Dunfield

On Geometric Aspects of Diffuse Groups  873–915

Skip Garibaldi and Holger P. Petersson Outer Automorphisms of Algebraic Groups

and a Skolem-Noether Theorem for Albert Algebras  917–954

Julien Bichon Gerstenhaber-Schack and Hochschild Cohomologies

of Hopf Algebras  955–986

Dominic Enders A Characterization of Semiprojectivity

for Subhomogeneous C^{∗}-Algebras  987–1049